a chain of mass m and length l is placed on a table with one-sixth of it hanging freely from the table edge .the amount of work done to pull the chain on the table is?

Grade:12th pass

1 Answers

Arun
5 years ago
Dear student
 
First of all calculate linear mass density of chain that is equal to mass/length. next since 1/6 length is hanging the centre of mass will rise to a height of 1/12length on pulling the chain up.next work done =m*g*h=(1/6L*m/L)*g*1/12L=1/72 mgl.
 
Regards
Arun